OpenStudy (anonymous):

The tangent of an acute angle of a right triangle is always less than one. This is false because say for example the two legs were 2 and 1 then if you were to tangent the angle opposite 2 the tangent would equal two. However, cos and sin will always be less than one, right?

OpenStudy (anonymous):

Right. \[\cos=\frac{ opp }{ hyp }~~~~and~~~\sin=\frac{ adj }{ hyp }\]hypotenuse is always going to be larger than the adjacent or opposite lengths. so your bottom number will always be larger than your top number

OpenStudy (phi):

*** , cos and sin will always be less than one ? *** always... the max value of sin or cos is 1 see plots: http://www.mathsisfun.com/algebra/trig-sin-cos-tan-graphs.html

OpenStudy (anonymous):

Yeah, that's what I was thinking to. It could equal one. But if you have a right triangle could it be 1? Or if it equals one is it now a line instead of a right triangle? @phi

OpenStudy (phi):

sin(x)= 1 means the angle is 90 degrees and the "right" triangle is no longer a triangle... it is just a line going straight up